There are 4 male students to every 8 female students at a local school. Select the statement showing the condition when there are 21 male students.
[A] 4/8=21/x [B]21/4=8/x [C] 4/8=x/21 [D] none of these

Answer by stanbon(75887)

There are 4 male students to every 8 female students at a local school. Select the statement showing the condition when there are 21 male students.
Use a proportion::
x/21 = 8/4
OR
4/8 = 21/x
===================
Cheers,
Stan H.
===================
[A] 4/8=21/x [B]21/4=8/x [C] 4/8=x/21 [D] none of these